# Robust Capacity Expansion Modelling for Renewable Energy Systems under Weather Uncertainty

### v3

@@ INTRODUCTION
leads to robust energy systems regardless of the initial time–series chosen. On average, this incurs additional cost of 7.87.8bn€ if no smoothing is performed. The average total cost of a robust system reaches 63.163.1bn€, with a range of [60.9,69.7][60.9,69.7]bn€. In comparison, using smoothing leads to slightly more expensive solutions. On average, making an energy system model robust incurs additional cost of 9.79.7bn€ ($17\%$). The average total cost of a robust system reaches 65.265.2bn€, with a range of [61.0,72.7][61.0,72.7]bn€.
Figure 5 gives the results of optimising each year independently for the smoothed MOD 1 / Demand increase. In comparison to non–smoothed MOD 1 / Demand increase, smoothed leads to on average lower investment increase in short–term battery storage ($25\%$ non–smoothed vs. $12\%$ for smoothed) and a significantly higher investment in CCGT ($86\%$ for non–smoothed vs. $36\%$ for smoothed). The higher costs may be due to the fact that additional artificial demand is added in time periods adjacent to those with previous supply gaps, which generates small supply demand gap time periods. The strong invest in CCGT compared to the non–smoothed modification suggests that CCGT power plants are used to offset those artificial demand gap time periods.
Figure 5: Total annual cost comparison from 198020191980-2019 for robust solutions using modification smoothed MOD 1 / Demand increase for the single node model in gurobipy, no temporal aggregation.
In summary, smoothed MOD 1 / Demand increase leads to feasible solutions, but care should be taken that artificial energy demands do not lead to excessive building of CCGT power plants. Convergence of MOD 1 / Demand increase can be slow, sometimes taking more than $20$ iterations for a single pair of years. This number of iterations was used as a cut-off criterion, as no improvements were observed after that during testing. This appears to be due to very small residual supply gaps of a few $GWh$ that get found and added to the model repeatedly. Given the small size of those supply gaps, the large overall production, and the fact that Gurobi  26 was used as a solver, which does not perform exact arithmetic, this may be caused by numerical instabilities. Using a suitable termination criterion (e.g. number of iterations or total supply gap less than some small number of GWh) counteracts this.
@@
is evaluated in Figures 6 and 7 for the model in gurobiy and the model in ETHOS.FINE, respectively, and Table 3, which give an overview of the results of generating robust energy system models for the five selected reference years in ETHOS.FINE.
On the one hand, applying MOD 2 / Synthetic time–series to the gurobipy model leads to robust solutions with the lowest cost increase on average. As can be seen in Figure 6, the total annual cost is similar across all $40$ years. On average, additional cost of 5.65.6bn€ or $10\%$ with a range of $[60.5,63.15]$ are the result of this modification. The overall deviation from the average robust solution is in the range $1\%,3\%]$ showing the effectiveness of the algorithm independent of the selected reference year. The consistently lower costs indicate that MOD 2 / Synthetic time–series is the best performing modification.
On the other hand, applying MOD 2 / Synthetic time–series in ETHOS.FINE is shown in Figure 7. After modifying the five selected reference years, the share of total annual cost for wind onshore decreases ($2\%-27\%$). Similarly, a decrease in total cost for transmission (electricity grid and hydrogen pipeline) is observed ($0\%-25\%$). A general increase in cost is seen for PV ($4\%-+15\%$), Li–ion batteries ($1\%-+26\%$) as well as the hydrogen sector ($9\%-+18\%$) for the robust energy system designs. The increase of PV can be explained by its below average, but still relevant, availability during dark lulls combined with Li–ion batteries to cover daily fluctuations. As visible in Figure 8, PV is mainly utilised together with Li–ion batteries to cover the fluctuating part of the electricity demand, the CCGT cover the bulk of the electricity demand, while the generation from wind is negligible. Hydrogen is utilised for electricity generation to a higher degree, since it can provide flexible additional energy supply, especially during dark lulls. The overall increase in cost compared to the average cost of each of the five reference years is $12\%-13\%$, compared to the weather year $2018$ it is $10\%-12\%$ and compared to the most expensive single year, which is a lower (dual) bound on the objective, it is $2.9\%-5\%$. Thus, MOD 2 / Synthetic time–series leads to robust and on average cheaper solutions in ETHOS.FINE similar to the results in the gurobipy model.
@@
Figure 6: Total annual cost comparison for optimising for from 198020191980--2019 for robust solutions using modification MOD 2 / Synthetic time–series gurobipy for the singe node model in gurobipy, no temporal aggregation.
Figure 7: Total annual cost comparison for optimising for 55 selected reference years for Germany. Columns marked with a * indicate the robust system designed via MOD 2 / Synthetic time–series ETHOS.FINE.
@@
Compared to the results of this work, Ryberg 12 estimates a residual load of about 6161GW and additional backup capacity required of about 2525GW for Germany. The difference to the 118118GW–125125GW found in this study can be attributed to the fact that in the integrated European setting that Ryberg 12 used, dark lulls can be partly suppressed by electricity transmitted from regions not hit by that dark lull as well as differences in demand data.
In either case, the cold dark lull period is the most critical for CCGT – their installed capacity is mainly driven by a single dark lull period, as shown in Figure 8. Figure 8 also illustrates the effects of MOD 2 / Synthetic time–series. The left graphic shows the result of the energy system model optimised for 1987 when testing its feasibility in 1994 reveals a supply gap. After applying MOD 2 / Synthetic time–series, the time period gets integrated into the optimisation problem and after reoptimising the supply can now be covered using existing capacities.
@@
The total costs average out to 66.066.0bn€ per year, with a range of $[61.2,72.1]$, which is slightly more than MOD 1 / Demand increase. Thus, in terms of costs, the latter is preferable. Figure 9 gives the results of optimising each year independently for MOD3 / Combine. Notably, the results for each year are very similar to each other, suggesting robust solutions share some traits. Cheap solutions contain less onshore wind.
@@
Figure 9: Total annual cost comparison for robust solutions based on single years using Modification MOD3 / Combine, a single node gurobipy model.
Capacity changes in robust solutions
@@
Compared to optimising individual years, cheap solutions to the robust model systematically use less onshore wind. This is plausible, since most years contain no extended (dark) lull periods coinciding with peak demands. In years without extended dark lulls, wind power provides stable and cheap energy, compared to PV that might require more storage and conversion units. However, in years with dark lulls this advantage disappears. As such, if costs for storage and conversion are priced in, optimising a year without a dark lull may lead to more investment in onshore wind than would be efficient. Integrating appropriate dark lull periods, as suggested in this work, might help counteract that effect, leading to a more balanced energy mix. A functioning capacity market, especially for backup technologies, is essential to ensure the needed capacities are installed and ready to generate electricity during dark lull periods. Overall, robust solutions were only $2-3\%$ more expensive compared to the lower (dual) bound given by the most expensive single year. Contrarily, a model based on average or recommended reference years systematically underestimates costs by over $10\%$.
@@
Lower investment in onshore wind capacity was a reoccurring pattern in the effects of modifications. The left graphic in Figure 10 shows the annual full load hours (FLH) for wind on– and offshore as well as PV compared to total annual costs for the respective models. For wind on– and offshore these are strongly correlated (Pearson correlation coefficients of 0.77-0.77 and 0.81-0.81 for wind on– and offshore, respectively). The FLH of PV and the TAC are nearly uncorrelated (Pearson correlation coefficient of 0.04-0.04). For wind this mirrors earlier results of Gotske et al. 23, who showed similar correlations for an European System.
Figure 10: The left graphic shows wind on– and offshore as well as PV full load hours compared to total annual cost for all years. The right graphic shows combined cost for wind and for PV as share of the total annual cost (TAC) compared to total annual cost for all years. Each dot represents one year. Lines indicate regression lines of the different energy generator types.
The right graphic in Figure 10 shows the share of total annual cost of PV and wind, combined for on– and offshore, compared to to the total annual cost. For wind these are negatively correlated (Pearson correlation coefficient of 0.59-0.59) while for PV these are strongly positively correlated (Pearson correlation coefficient of $0.59$) indicating that as full load hours of wind drop, wind capacity gets replaced with PV capacity. The two diagrams in Figure 10 together show that the total annual cost depends strongly on the availability of wind. In weather years with low full load hours for wind, an increase in PV capacities is observed indicating a higher reliance on PV in general which is also reflected in the robust years as investment increases for PV capacities for MOD 2 / Synthetic time–series and MOD3 / Combine.
@@
The approaches outlined in this work can be used to make any energy system model more robust to uncertain time–series data. As we showed, this serves to makes models more realistic and more suitable for real–world application.
Since uncertainties in demands or costs, which appear in the right–hand side (demands) or the objective (cost) of a MIP can be reformulated as constraint–wise uncertainty, the methods outlined in this work can be applied to cost uncertainty as well. Another natural usage for the modifications proposed in this work is as part of a Benders decomposition framework. However, this was not the focus of this work, but the constraints that are added in each modification can equally well be used as feasibility/optimality cuts, as they serve to invalidate significant parts of the solution space. Here, the algorithmic performance for achieving robustness can be seen as a proxy for their potential value as Benders cuts.
@@
In this case study, only one time–series for electricity demand in $2050$ was used. The availability of future demand data is limited and is difficult to compare if taken from different sources. The selected data from Forschungsstelle für Energiewirtschaft e. V. (FfE) 27, 28, 29, 30 was chosen since it includes a severe cold period rarely observed in Germany. Therefore a high degree of robustness can be assured. Further demand time–series might still provide additional insights. This is especially the case if using weather years with an overall cold winter time and therefore a high amount of heating degree days, but also varying demand distribution and even lower temperatures.
Assuming that years with disadvantageous distribution of sunny and wind hours as well as low full load hours are the exception, allowing the last time step to have a lower state of charge than the first would lower the conservatism of the system and reduce cost. This would require a measure of robustness that fully protects against a certain base uncertainty set, but allows using up some stored hydrogen for outlier events. One possible approach for that is outlined in Bärmann et al. 31. At the same time, the model only provides an operational schedule under the assumption of perfect foresight within one year. In a more complex model setting, computing an operational schedule dynamically throughout the year might lead to some efficiency losses in the usage of energy, requiring additional investment to counteract this. Lastly, the algorithmic approach proposed in this work is not designed to address long–term effects of climate change, but can be utilised for time–series data that include such effects just the same.
Methods
@@
After solving a initial CAPEX problem, the proposed algorithm consists of four main steps, as shown in Figure 1: 1. feasibility testing using data of other weather years  [^1] In robust optimisation, scenarios are commonly used as a term to denote different possible realisations of uncertainty. In this setting, each scenario consists of one year–long time–series of weather data for each type of renewable energy affected by the weather. As such, this is a discrete uncertainty set. While we introduce these notions for weather patterns only, the theory applies to any finite scenario set and their convex combination $\mathcal{U}=\textrm{conv}(u_{1},...u_{n})$. , 2. assessing loss of load/supply gaps, 3. identification of reasons for infeasibility, specifically critical time periods to derive modifications and, 4. solving the modified CAPEX problem before repeating steps 1-4. Before describing these steps, we define robustness for energy system models.
@@
As noted before, general ARO has a bilevel structure: first stage decisions are made, some (worst–case) uncertainty realises, but then second stage decisions can be then made based on the then known uncertainty and the fixed first stage decisions. For energy system modelling, we let the capacity decision variables be the first stage, since new power plants can not be build ad hoc to react to weather variations. The second stage decisions consist of operational decisions that can be made based on known weather variations. Thus, all second stage costs are operational expenditures (OPEX). We assume that the OPEX can be given as a flat percentage of investment costs/capital expenditure ($c_{CAPEX}^{\top}$), which is a reasonable assumption for many renewables such as wind power and solar photovoltaics 32.
Via those simplifications, we remove costs from the second stage decisions. Mathematically, this is a crucial simplification as it allows us to collapse the ”min-max-min” objective of general ARO into a simple minimisation problem. A second motivation for doing that is that our approach emulates the principle of a Bender’s decomposition (see the most recent book by Hooker 33 for an overview of Bender’s decomposition). By removing the second stage costs, we remove the need for optimality cuts, which significantly simplifies computations. This results in the following, simplified ARO model:
@@
Table 4: Energy system components considered for development of the proposed methodology for optimising energy system models for Germany.
Appendix A gives an overview of the simplified model used for implementing constraints and variable demands to allow for an easy and fast testing of the multiple solution modifications used below. The full code, model and data are publicly available via GitHub 34.
@@
A solution to the mixed–integer program ARO-ESM that is optimal and feasible for one year can lead to supply gaps for other weather years. We call an energy system models robust against a set of weather years $\mathcal{U}$, if and if only it can be operated without supply gaps for any weather year $u\in\mathcal{U}$:
@@
Remark that this definition explicitly defines robustness relative to a known uncertainty set $\mathcal{U}$. Furthermore, it requires non–decreasing storage, which implies long–term supply security. That means even multiple sequential years with low full load hours for renewables will not deplete storage levels. This is a risk–averse strategy, as increased energy supply in beneficial weather years can not be used to offset lack of supply in bad years.
@@
Here, $\delta$ is a vector of supply gaps. It is implied that $\delta_{ij}$ is only added to constraints that cover energy supply/demand. Note that in the mixed integer program above, $x_{u_{i}}$, the capacity expansion, is a fixed input parameter retrieved from the selected year. Now an energy system designed for weather year $u_{i}$ is robust for any weather year $u_{j}$ if and if only there exists an solution to the mixed-integer program $COMP_{j}^{i}$ above such that $\delta=0$. If not, $\delta_{ij}$ encodes where supply gaps are and how large they are. [^2] Note that this is a relation between two $x_{s_{i}},x_{s_{j}}$ solutions, not their respective underlying time–series data, since there may be multiple (near-)optimal solutions for a given weather and demand year. Computationally, this never proved to be an issue, so these notions are used interchangeably throughout this work. However, care should be taken if multiple capacity decisions have equal pricing.

### v4

@@ A Novel Algorithm for Computing Robust ESMs
leads to robust energy systems regardless of the initial time–series chosen. On average, this incurs additional cost of 7.87.8bn€ if no smoothing is performed. The average total cost of a robust system reaches 63.163.1bn€, with a range of [60.9,69.7][60.9,69.7]bn€. In comparison, using smoothing leads to slightly more expensive solutions. On average, making an ESM robust incurs additional cost of 9.79.7bn€ ($17\%$). The average total cost of a robust system reaches 65.265.2bn€, with a range of [61.0,72.7][61.0,72.7]bn€.
Figure 6 gives the results of optimising each year independently for the smoothed MOD 1 / Demand increase. While on average more expansive than non–smoothed MOD 1 / Demand increase, the smoothed algorithm leads to on average lower investment increase in short–term battery storage ($25\%$ non–smoothed vs. $12\%$ for smoothed) and a significantly higher investment in CCGT ($86\%$ for non–smoothed vs. $36\%$ for smoothed). The higher costs may be due to the fact that additional artificial demand is added in time periods adjacent to those with previous supply gaps, which generates small supply demand gap time periods. The strong invest in CCGT compared to the non–smoothed modification suggests that CCGT power plants are used to offset those artificial demand gap time periods.
Figure 6: TAC comparison from 198020191980-2019 for robust solutions using modification smoothed MOD 1 / Demand increase for the single node model in gurobipy, no temporal aggregation.
In summary, smoothed MOD 1 / Demand increase leads to feasible solutions, but care should be taken that artificial energy demands do not lead to excessive building of CCGT power plants. Convergence of MOD 1 / Demand increase can be slow, sometimes taking more than $20$ iterations for a single pair of years. This number of iterations was used as a cut-off criterion, as no improvements were observed after that during testing. This appears to be due to very small residual supply gaps of a few $GWh$ that get found and added to the model repeatedly. Given the small size of those supply gaps, the large overall production, and the fact that Gurobi  24 was used as a solver, which does not perform exact arithmetic, this may be caused by numerical instabilities. Using a suitable termination criterion (e.g. number of iterations or total supply gap less than some small number of GWh) counteracts this.
@@
is evaluated in Figures 7 and 8 for the model in gurobipy and the model in ETHOS.FINE, respectively, and Table 2, which give an overview of the results of generating robust ESMs. In the ETHOS.FINE model the five most expensive solutions based on the single years $1987,1996,2009,2010$ and $2013$ were selected of the $40$ energy system designs to make them robust using MOD 2 / Synthetic time–series. These are hereon referred to as the five reference years.
On the one hand, applying MOD 2 / Synthetic time–series to the gurobipy model leads to robust solutions with the lowest cost increase on average. As can be seen in Figure 7, the TAC is similar across all $40$ years. On average, additional cost of 5.65.6bn€ or $10\%$ with a range of $[60.5,63.15]$ are the result of this modification. The overall deviation from the average robust solution is in the range $1\%,3\%]$ showing the effectiveness of the algorithm independent of the selected reference year. The consistently lower costs indicate that MOD 2 / Synthetic time–series is the best performing modification.
On the other hand, applying MOD 2 / Synthetic time–series in ETHOS.FINE is shown in Figure 8. After modifying the five selected reference years with MOD 2 / Synthetic time–series to make the solutions robust, the share of TAC for wind onshore decreases ($2\%-27\%$). Similarly, a decrease in total investment in transmission (electricity grid and hydrogen pipeline) is observed ($0\%-25\%$). A general increase in investment is seen for PV ($4\%-+15\%$), Li–ion batteries ($1\%-+26\%$) as well as the hydrogen sector ($9\%-+18\%$) for the robust energy system designs. The increase of PV can be explained by its below average, but still relevant, availability during dark lulls combined with Li–ion batteries to cover daily fluctuations. As visible in Figure 9, PV is mainly utilised together with Li–ion batteries to cover the fluctuating part of the electricity demand, the CCGT cover the bulk of the electricity demand, while the generation from wind is negligible. Hydrogen is utilised for electricity generation to a higher degree, since it can provide flexible additional energy supply, especially during dark lulls. The overall increase in cost compared to the average cost of each of the five reference years is $12\%-13\%$, compared to the weather year $2018$ it is $10\%-12\%$ and compared to the most expensive single year, which is a lower (dual) bound on the objective, it is $2.9\%-5\%$. Thus, MOD 2 / Synthetic time–series leads to robust and on average cheaper solutions in ETHOS.FINE similar to the results in the gurobipy model.
@@
Figure 7: TAC comparison for optimising for from 198020191980-2019 for robust solutions using modification MOD 2 / Synthetic time–series gurobipy for the singe node model in gurobipy, no temporal aggregation.
Figure 8: TAC comparison for optimising for 55 selected reference years for Germany. Columns marked with a * indicate the robust system designed via MOD 2 / Synthetic time–series ETHOS.FINE.
@@
Compared to the results of this work, Ryberg 48 estimates a residual load of about 6161GW and additional backup capacity required of about 2525GW for Germany. The difference to the 118118GW–125125GW found in this study can be attributed to the fact that in the integrated European setting that Ryberg 48 used, dark lulls can be partly suppressed by electricity transmitted from regions not hit by that dark lull as well as differences in demand data.
In either case, the cold dark lull period is the most critical for CCGT – their installed capacity is mainly driven by a single dark lull period, as shown in Figure 9. Figure 9 also illustrates the effects of MOD 2 / Synthetic time–series. The left graphic shows the result of the ESM optimised for 1987 when testing its feasibility in 1994 reveals a supply gap. After applying MOD 2 / Synthetic time–series, the time period gets integrated into the optimisation problem and after reoptimising the supply can now be covered using existing capacities.
@@
The total costs average out to 66.066.0bn€ per year, with a range of $[61.2,72.1]$, which is slightly more than MOD 1 / Demand increase. Thus, in terms of costs, the MOD 1 / Demand increase is preferable. Figure 10 gives the results of optimising each year independently for MOD 3 / Combine. Notably, the results for each year are very similar to each other, suggesting robust solutions share some traits. Cheap solutions contain less onshore wind.
@@
Figure 10: TAC comparison for robust solutions based on single years using MOD 3 / Combine, a single node gurobipy model.
Comparison of Modifications
All three featured approaches generate robust solutions, i.e. solutions with no supply gaps, subject to numerical tolerances. The modifications’ performance is summarised in Table 3.
Modification Increased Investment Average Least
MOD 1 / Demand increase Li-Ion, $H_{2}$-CCGT, Salt caverns +5.7% +2.0%
Smoothed MOD 1 / Demand increase $H_{2}$-CCGT, Salt caverns +9.2% +2.4%
MOD 2 / Synthetic time–series gurobipy PV, $H_{2}$-CCGT, Salt caverns, $H_{2}$-Electrolysers +2.5% +1.6%
MOD 2 / Synthetic time–series ETHOS.FINE PV, $H_{2}$-CCGT, Salt caverns, $H_{2}$-Electrolysers +3.7% +2.9%
MOD 3 / Combine PV, Wind, $H_{2}$-CCGT, Salt caverns, $H_{2}$-Electrolysers +10.6% +2.5%
Table 3: Comparison of the different modifications in terms of convergence and performance. Average and least cost increase compared to best lower (dual) bound, the highest TAC of a single unmodified year.
The results show a moderate increase in investment for robust models leading to TAC increases of $1.6-2.9\%$ compared to the best available lower (dual) bound on the cost of a robust solution. The dual bound is given by the maximum total cost for a single year, since the cost of a robust solution cannot be lower than the cost of a single year. Thus, the highest cost year is also the only one that could be robust [^1] Clearly, this can be shown via a proof by contradiction, since each solution is assumed to be optimal and thus there must not be a feasible solution with a lower objective value.. For gurobipy, the lower bound is 59.759.7bn€ annually. For the model in, ETHOS.FINE it is 113.6113.6bn€ annually.
A upper (primal) bound is given by the maximum capacity for each technology per region and year, since this will be feasible for all years. This results in a range of [113.6,195.66][113.6,195.66]bn€ annually for an optimal robust energy system in ETHOS.FINE. For the gurobipy model, the range for optimal robust solutions is [59.7,67.6][59.7,67.6]bn€ annually.
Capacity changes in near-optimal robust solutions
It is important to acknowledge that any single robust solution may not accurately capture the properties of robust solutions as a whole, especially if considering the space of near-optimal solutions that may be relevant to decision makers, see the recommendations by Lombardi et al. 36. While we did not explicitly perform modelling-to-generate-alternatives on the solution space, computing $40$ extremal solutions on single years and iteratively robustifying them gives us a sample of $40$ near-optimal robust solutions that allow us to infer some properties of near-optimal robust solutions:
@@
Compared to optimising individual years, near-optimal solutions to the robust model systematically use less onshore wind. This is plausible, since most years contain no extended (dark) lull periods coinciding with peak demands. In years without extended dark lulls, wind power provides stable and cheap energy, compared to PV that might require more storage and conversion units. However, in years with dark lulls this advantage disappears. As such, if costs for storage and conversion are priced in, optimising a year without a dark lull may lead to more investment in onshore wind than would be efficient. Integrating appropriate dark lull periods, as suggested in this work, might help counteract that effect, leading to a more balanced energy mix. A functioning capacity market, especially for backup technologies, is essential to ensure the needed capacities are installed and ready to generate electricity during dark lull periods. Overall, robust solutions were only $2-3\%$ more expensive compared to the lower (dual) bound given by the most expensive single year. Contrarily, a model based on average or recommended reference years systematically underestimates costs by over $10\%$.
Convergence
The gurobipy implementations of MOD 1 / Demand increase and MOD 3 / Combine tended to quickly converge. However, sometimes residual load differences in the order of MWh/kWh remained. In these cases, the algorithm was terminated after $20$ iterations. MOD 2 / Synthetic time–series needed between one and $54$ iterations to converge in the gurobipy model. The number of iterations generally decreased for higher annual cost of individual years, but was mainly due to overwriting critical time periods where priority was given to critical time periods selected in earlier iteration steps. In ETHOS.FINE this issue was prevented by evaluating critical time periods on the modified data largely preventing overwriting. The calculations then needed between one and eight iterations for the five reference years. Here, one iteration incurs the same computational load of solving one ESM, or less if warm–starts help reducing computation times.
@@
Lower investment in onshore wind capacity was a reoccurring pattern in the effects of modifications. The left graphic in Figure 11 shows the annual full load hours (FLH) for wind on– and offshore as well as PV compared to TAC for the respective models. For wind on– and offshore these are strongly correlated (Pearson correlation coefficients of 0.77-0.77 and 0.81-0.81 for wind on– and offshore, respectively). The FLH of PV and the TAC are nearly uncorrelated (Pearson correlation coefficient of 0.04-0.04). For wind this mirrors earlier results of Gotske et al. gøtske2024designingsectorcoupledeuropeanenergy, who showed similar correlations for an European System.
Figure 11: The left graphic shows wind on– and offshore as well as PV full load hours compared to TAC in for all years. The right graphic shows combined cost for wind and for PV as share of the TAC compared to TAC for all years. Each dot represents one year. Lines indicate regression lines of the different energy generator types.
The right graphic in Figure 11 shows the share of TAC of PV and wind, combined for on– and offshore, compared to to the TAC. For wind these are negatively correlated (Pearson correlation coefficient of 0.59-0.59) while for PV these are strongly positively correlated (Pearson correlation coefficient of $0.59$) indicating that as full load hours of wind drop, wind capacity gets replaced with PV capacity. The two diagrams in Figure 11 together show that the TACs depend strongly on the availability of wind. In weather years with low full load hours for wind, an increase in PV capacities is observed indicating a higher reliance on PV in general which is also reflected in the robust years as investment increases for PV capacities for MOD 2 / Synthetic time–series and MOD 3 / Combine.
@@
[…]
Since uncertainties in demands or costs, which appear in the right–hand side (demands) or the objective (cost) of a mixed-integer linear program (MILP) can be reformulated as constraint–wise uncertainty, the methods outlined in this work can be applied to cost uncertainty as well. Another natural usage for the modifications proposed in this work is as part of a Benders decomposition framework. However, this was not the focus of this work, but the constraints that are added in each modification can equally well be used as feasibility/optimality cuts, as they serve to invalidate significant parts of the solution space. Here, the algorithmic performance for achieving robustness can be seen as a proxy for their potential value as Benders cuts.
@@
The main drawback in this case study is that only one time–series for electricity demand in $2050$ was used. The availability of future demand data is limited and is difficult to compare if taken from different sources. The selected data from Forschungsstelle für Energiewirtschaft e. V. (FfE) 17, 15, 18, 16 was chosen since it includes a severe cold period rarely observed in Germany. In combination with the weather years the resulting operational conditions revealed several periods pairing low electricity generation with high demand as can be seen in Figure 5. Therefore a high degree of robustness can be assured. Further demand time–series might still provide additional insights. This is especially the case if using weather years with an overall cold winter time and therefore a high amount of heating degree days, and thus varying total and local demand. The weather data used are based on historical weather years leaving out the effects of climate change, which very likely will influence the design of energy systems in the future and should be included in studies using this methodologies. The model used to develop the presented method is a Germany model without the possibility of imports or exports. While this island model approach was necessary for method development, the interconnectedness of the European energy system, as well as (cold) dark lulls covering large parts of Europe, is likely crucial, but has been neglected in this study. The Germany model proved sufficient for model development, but a more general model is necessary for more holistic recommendations. Assuming that years with disadvantageous distribution of sunny and wind hours as well as low full load hours are the exception, allowing the last time step to have a lower state of charge than the first would lower the conservatism of the system and reduce cost. This would require a measure of robustness that fully protects against a certain base uncertainty set, but allows using up some stored hydrogen for outlier events. One possible approach for that is outlined in Bärmann et al. 2. At the same time, the model only provides an operational schedule under the assumption of perfect foresight within one year. In a more complex model setting, computing an operational schedule dynamically throughout the year might lead to some efficiency losses in the usage of energy, requiring additional investment to counteract this. Perfect foresight is assumed for the optimization of energy systems. This provides the advantage of being able to plan capacities with full information about cold dark lulls and overall energy generation. In a real–world setting a higher degree of conservatism leading to additional generation or storage would be needed.
Summary of Implications for Planning Energy Systems
Our work contributes to the mounting evidence that suggest that using a fixed reference year and planning a renewable energy system based on that is unsuitable for practice 50, 45, 39, 25, 46, 22. We show that not only does this lead to misallocation of resources, but to a systematic underestimation of costs for and investment in storage and conversion, i.e. Li-Ion battery capacity and CCGT. Thus, if energy systems are planned based on single years, policy makers need to separately assess how much, not whether, additional storage and conversion capacities are needed and be aware of the fact that this will incur additional costs. For practitioners who model energy systems, we propose three workable approaches that can ensure energy system designs are robust against a range of weather realisations. We also note that the total amount of wind hours per year is strongly correlated with energy system costs. Thus, for a conservative estimate, low-wind reference years are better suited.
STAR Methods
Key Resource Table
REAGENT or RESOURCE SOURCE
Deposited data
Load Curves of the Tertiary Sector – eXtremOS solidEU Scenario (Europe NUTS-3) opendata.ffe.de / tertiary load curves
Load Curves of the Household Sector – eXtremOS solidEU Scenario (Europe NUTS-3) opendata.ffe.de / household load curves
Load Curves of the Transport Sector – eXtremOS solidEU Scenario (Europe NUTS-3) opendata.ffe.de / transport load curves
Load Curves of the Industry Sector – eXtremOS solidEU Scenario (Europe NUTS-3) opendata.ffe.de / industry load curves
40 years of wind and PV data in hourly resolution renewables.ninja / wind PV data
Tool for Renewable Energy Potentials – Database zenodo.org / renewable potentials
Council Regulation (EU) No 1059/2003 eur-lex.europa.eu / EU regulation 1059
Software and algorithms
Full code, model and data github.com / robust capacity expansion
Method Details
@@
After solving a initial CAPEX problem, the proposed algorithm consists of four main steps, as shown in Figure 1: 1. feasibility testing using data of other years  [^2] In robust optimisation, scenarios are commonly used as a term to denote different possible realisations of uncertainty. In this setting, each scenario consists of one year–long time–series of data for each type of renewable energy in combination with the demand affected by the weather. As such, this is a discrete uncertainty set. While we introduce these notions for weather patterns only, the theory applies to any finite scenario set and their convex combination $\mathcal{U}=\textrm{conv}(u_{1},...u_{n})$., 2. assessing loss of load/supply gaps, 3. identification of reasons for infeasibility, specifically critical time periods to derive modifications and, 4. solving the modified CAPEX problem before repeating steps 1-4. Before describing these steps, we define robustness for ESMs.
@@
As noted before, general ARO has a bilevel structure: first stage decisions are made, some (worst–case) uncertainty realises, but then second stage decisions can be then made based on the then known uncertainty and the fixed first stage decisions. For energy system modelling, we let the capacity decision variables be the first stage, since new power plants can not be build ad hoc to react to weather variations. The second stage decisions consist of operational decisions that can be made based on known weather variations. Thus, all second stage costs are operational expenditures (OPEX). We assume that the OPEX can be given as a flat percentage of investment costs/capital expenditure ($c_{CAPEX}^{\top}$), which is a reasonable assumption for many renewables such as wind power and solar PV 30. This is assumed for all technologies included, but the fuel cost for hydrogen, which are endogenous, since hydrogen gas turbines can only use hydrogen produced by the system itself.
Via those simplifications, we remove costs from the second stage decisions. Mathematically, this is a crucial simplification as it allows us to collapse the ”min-max-min” objective of general ARO into a simple minimisation problem. A second motivation for doing that is that our approach emulates the principle of a Bender’s decomposition (see the most recent book by Hooker 29 for an overview of Bender’s decomposition). By removing the second stage costs, we remove the need for optimality cuts, which significantly simplifies computations. This results in the following, simplified ARO model:
@@
Table 4: Energy system components considered for development of the proposed methodology for optimising ESMs for Germany.
Appendix A gives an overview of the simplified model used for implementing constraints and variable demands to allow for an easy and fast testing of the multiple solution modifications used below. The full code, model and data are publicly available via GitHub 33.
@@
A solution to the mixed–integer program ARO-ESM that is optimal and feasible for one year can lead to supply gaps for other years. Furthermore, since MOD 2 / Synthetic time–series modifies the weather data resulting in artificial weather years, which are not based on atmospheric physics. The feasibility testing serves as a method to ensure that optimized energy systems can be operated in historic weather years. Therefore, feasibility testing has two purposes: First, to detect supply gaps. Second, to ensure that optimized energy systems are realistic. We call an ESM robust against a set of years $\mathcal{U}$, if and if only it can be operated without supply gaps for any year $u\in\mathcal{U}$:
@@
Remark that this definition explicitly defines robustness relative to a known uncertainty set $\mathcal{U}$. Furthermore, it requires non–decreasing storage, which implies long–term supply security. That means even multiple sequential years with low full load hours for renewables will not deplete storage levels. This is a risk–averse strategy, as increased energy supply in beneficial years can not be used to offset lack of supply in bad years.
@@
Here, $\delta$ is a vector of supply gaps. It is implied that $\delta_{ij}$ is only added to constraints that cover energy supply/demand. Note that in the mixed integer program above, $x_{u_{i}}$, the capacity expansion, is a fixed input parameter retrieved from the selected year. Now an energy system designed for year $u_{i}$ is robust for any year $u_{j}$ if and if only there exists an solution to the MILP $COMP_{j}^{i}$ above such that $\delta=0$. If not, $\delta_{ij}$ encodes where supply gaps are and how large they are. [^3] Note that this is a relation between two $x_{s_{i}},x_{s_{j}}$ solutions, not their respective underlying time–series data, since there may be multiple (near-)optimal solutions for a given year. Computationally, this never proved to be an issue, so these notions are used interchangeably throughout this work. However, care should be taken if multiple capacity decisions have equal pricing.